Vaughn Company applies manufacturing overhead to jobs on the basis of machine hours used. Overhead costs are estimated to total $283,725 for the year, and machine usage is estimated at 126,100 hours.

For the year, $303,850 of overhead costs are incurred and 130,200 hours are used.

Compute the manufacturing overhead rate for the year. (Round answer to 2 decimal places, e.g. 1.25.)
Manufacturing overhead rate $ per machine hour

SHOW LIST OF ACCOUNTS

What is the amount of under- or overapplied overhead at December 31?
Manufacturing Overhead $

OverappliedUnderapplied

SHOW LIST OF ACCOUNTS

Prepare the adjusting entry to assign the under- or overapplied overhead for the year to cost of goods sold. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit
Answer #1

Manufacturing Overhead rate

= Estimated overhead/Estimated machine hours

= 283,725/126,100

= 2.25 per machine hour

Applied overhead = 2.25*130,200 = 292,950

Underapplied overhead = 303,850 - 292,950

Underapplied overhead = 10,900

General journal Debit Credit
Cost of Goods Sold 10900
Manufacturing overhead 10900

(to record underapplied overhead)